The sum of a number,1/6 of that number,2 1/2, and 7 is 12 1/2. Find the number
Answer by fractalier(6550)

We have
x + (1/6)x + 5/2 + 7 = 25/2
Now simplifying we get
(7/6)x + 19/2 = 25/2
Subtract 19/2 from both sides
(7/6)x = 6/2 = 3
Now multiply both sides by 6/7 and get
x = 18/7